4. The current stock price of ABC, Corp. is $300 per share. ABC, Corp. just paid an annual per share dividend of $3. Over the past 5 years, this dividend has been grown annually at a rate of 35%. A respected analyst assumes that the current growth rate of dividends will hold up for the next 5 years, after which dividend growth will slow to 8% annually. Use the solver in Excel to compute the cost of equity?

A. 32.76% B. 12.88% C. 11.52% D. 11.15%
Answer #1

D. 11.15%

C E F G H I J K L B cost of equity 0.1 (initial guess) Solver Parameters year 1 Set Objective: SDS12 Dividend Present value 4
cost of equity 0.111477 11.15% year Dividend Present value 4.05 3.643800756 2 5.4675 4.425761318 3 7.381125 5.375530812 4 9.9 А cost of equity 0.111476798730908 =C1 year m Dividend =3*1.354B4 =3*1.35^B5 =3*1.354B6 =3*1.35^B7 =3*1.35^B8 Present value =

Please do rate me and mention doubts, if any, in the comments section.